# Unleashing Semantic and Geometric Priors for 3D Scene Completion

### v1

@@ Abstract
Camera-based 3D semantic scene completion (SSC) provides dense geometric and semantic perception for autonomous driving and robotic navigation. However, existing methods rely on a coupled encoder to deliver both semantic and geometric priors, which forces the model to make a trade-off between conflicting demands and limits its overall performance. To tackle these challenges, we propose FoundationSSC, a novel framework that performs dual decoupling at both the source and pathway levels. At the source level, we introduce a foundation encoder that provides rich semantic feature priors for the semantic branch and high-fidelity stereo cost volumes for the geometric branch. At the pathway level, these priors are refined through specialised, decoupled pathways, yielding superior semantic context and depth distributions. Our dual-decoupling design produces disentangled and refined inputs, which are then utilised by a hybrid view transformation to generate complementary 3D features. Additionally, we introduce a novel Axis-Aware Fusion (AAF) module that addresses the often-overlooked challenge of fusing these features by anisotropically merging them into a unified representation. Extensive experiments demonstrate the advantages of FoundationSSC, achieving simultaneous improvements in both semantic and geometric metrics, surpassing prior bests by +0.23 mIoU and +2.03 IoU on SemanticKITTI. Additionally, we achieve state-of-the-art performance on SSCBench-KITTI-360, with 21.78 mIoU and 48.61 IoU. The code will be released upon acceptance.
@@ Introduction
Semantic Scene Completion (SSC) infers the complete geometry and semantics of a 3D scene from partial observations. It is essential for high-level 3D perception applications, such as intelligent driving (Zheng et al. 2024) and robotic navigation (Wu et al. 2025). Early SSC methods primarily utilised LiDAR input (Song et al. 2017; Roldao, De Charette, and Verroust-Blondet 2020), but they faced challenges due to high hardware costs and limited resolution. As a result, there has been a recent shift towards camera-based SSC methods, which are gaining significant attention for their lower costs and richer texture information.
@@
In this paper, we propose FoundationSSC, a novel SSC framework designed to unleash the semantic and geometric priors from foundation models, as shown in Fig. 1 (b). Our approach operates on a core principle of dual decoupling. It begins with a unified foundation encoder that harnesses the power of VFMs to provide separate priors for the semantic and geometric branches, thereby addressing the trade-off at the feature source. These distinct priors are then processed through specialised, decoupled pathways aimed at maximising their potential. The resulting high-quality 2D representations are transformed into 3D space using a hybrid view transformation. Additionally, we introduce a novel Axis-Aware Fusion (AAF) module, which fuses the complementary 3D feature volumes in a way that respects their anisotropic characteristics, resulting in a unified 3D representation. Through these integrated designs, FoundationSSC effectively resolves the semantic-geometric coupling issues that have limited previous work, producing high-quality 2D and 3D feature representations. Consequently, our framework achieves state-of-the-art performance on SemanticKITTI (Behley et al. 2019) and SSCBench-KITTI-360 (Liao, Xie, and Geiger 2022) datasets, simultaneously enhancing both geometric (Intersection over Union, IoU) and semantic (mean IoU, mIoU) metrics. In summary, our contributions are as follows:
@@
This section provides a concise overview of recent developments in VFMs and methods for camera-based SSC.
@@
Our proposed method, FoundationSSC, is fundamentally different from these existing approaches. It not only utilises robust foundation priors but also introduces a novel dual-decoupling architecture to fully harness both semantic and geometric foundation priors simultaneously.
@@
MonoScene (Cao and De Charette 2022) is a pioneering effort that extends the SSC task to cameras by employing a Features Line of Sight Projection (FLoSP) for projecting 2D features into 3D space. TPVFormer (Huang et al. 2023) introduces an efficient tri-perceptive view representation, created from image features. VoxFormer (Li et al. 2023a) employs a sparse-to-dense strategy, which propagates image features to dense voxels through deformable attention mechanism. OccFormer (Zhang, Zhu, and Du 2023) builds upon the Lift-Splat-Shoot (LSS) paradigm to facilitate feature lifting using both local and global transformer pathways.
Subsequent works have proposed more specialised solutions, often leveraging temporal cues from multiple frames (Li et al. 2024a; Lee et al. 2025), utilising more accurate geometric information obtained from stereo matching (Li et al. 2024b; Xue et al. 2024; Yu et al. 2024), exploiting spatial biases (Wang and Tong 2024; Bae, Ha, and Kim 2025), and integrating richer semantic knowledge derived from instance-level features (Jiang et al. 2024) or vision-language models (Wang et al. 2025; Cui et al. 2025). However, despite the diversity of these approaches, they are often limited by a tightly coupled architecture, which results in a trade-off between semantic and geometric features.
To address this limitation, we propose FoundationSSC, which disentangles semantic and geometric features at both the source and processing pathway levels. This dual-decoupling approach enables simultaneous improvements in both geometric and semantic metrics, effectively resolving the trade-off that has constrained earlier work.
@@
  • Dense depth map $\mathbf{Z}$: The final disparity map predicted by FoundationStereo, which then converted to a metric depth map using camera intrinsics. This is primarily used for generating voxel proposals in the view transformation and geometric prior for the semantic context branch.
@@
Disparity-to-Depth Volume Mapping
@@
To restore spatial structure, $\mathbf{\tilde{V}}_{depth}$ are reshaped into 3D volumes $\mathbf{V}_{depth}\in\mathbb{R}^{D_{depth}\times H\times W}$. A refinement module then employ a shallow 3D CNN encoder $\phi(\cdot)$ to enhance spatial coherence and smoothness, followed by a softmax operation along the depth dimension to produce the probabilistic depth distribution, $\mathbf{D}\in\mathbb{R}^{D_{depth}\times H\times W}$:
@@
Hybrid View Transformation.
@@
Axis-Aware Fusion
@@
To train the FoundationSSC, we follow MonoScene (Cao and De Charette 2022), utilise affinity loss $\mathcal{L}_{scal}^{geo}$ and $\mathcal{L}_{scal}^{sem}$ to optimise the scene-wise and class-wise metric, in conjunction with the cross-entropy loss $\mathcal{L}_{ce}$ weighted by class frequencies. Additionally, we employ two auxiliary losses: a 2D semantic segmentation loss $\mathcal{L}_{seg}$ for feature enrichment and a depth loss $\mathcal{L}_{d}$ (Li et al. 2024e) for geometric supervision. The total loss is a weighted sum of these components:
@@
where $\lambda_{d}$ and $\lambda_{seg}$ are the weights for the auxiliary losses. Following preliminary experiments, we set $\lambda_{d}=0.001$ and $\lambda_{seg}=1$, respectively.
@@
Ablation Study for Foundation Encoder
@@
Ablation Study for DDVM
We validate our DDVM module against various depth generation strategies in Table 5. The CGFormer (Yu et al. 2024) depth refinement module, which fuses monocular distribution with a pre-processed stereo depth map using neighbour attention, achieves 19.82 mIoU and 47.87 IoU. A key limitation of such a method is its reliance on information-lossy depth maps, which discard valuable probabilistic cues from the original stereo matching. We then explore the direct use of the disparity cost volume. A naive approach using analytical resampling (AR) degrades performance to 19.59 mIoU, confirming that merely accessing the cost volume is insufficient. In contrast, our DDVM, by replacing rigid resampling with a learnable mapping, achieves a state-of-the-art mIoU of 20.36. The performance gain demonstrates DDVM’s superior ability to preserve and leverage fine-grained geometric cues for SSC.
@@
Ablation Study for AAF
Table 6 compares the AAF module against alternative 3D feature fusion strategies. We establish a baseline by using only the refined volume $\mathbf{F}_{vt}$, mirroring prior work, which yields 19.56 mIoU. A simple fusion approach using a vanilla 3D channel attention (3D-CA) module improves performance to 20.08 mIoU. However, the effectiveness of 3D-CA is constrained by an isotropic global pooling, a design that neglects directional cues critical for 3D understanding. In contrast, our AAF module is explicitly designed to be anisotropic. The module’s axis-specific fusion mechanism achieves the highest mIoU of 20.36. The substantial performance gain over the 3D-CA provides compelling evidence that modelling the anisotropic nature of 3D scenes is crucial for effective feature fusion.

### v2

@@ Abstract
Camera-based 3D semantic scene completion (SSC) provides dense geometric and semantic perception for autonomous driving and robotic navigation. However, existing methods rely on a coupled encoder to deliver both semantic and geometric priors, which forces the model to make a trade-off between conflicting demands and limits its overall performance. To tackle these challenges, we propose FoundationSSC, a novel framework that performs dual decoupling at both the source and pathway levels. At the source level, we introduce a foundation encoder that provides rich semantic feature priors for the semantic branch and high-fidelity stereo cost volumes for the geometric branch. At the pathway level, these priors are refined through specialised, decoupled pathways, yielding superior semantic context and depth distributions. Our dual-decoupling design produces disentangled and refined inputs, which are then utilised by a hybrid view transformation to generate complementary 3D features. Additionally, we introduce a novel Axis-Aware Fusion (AAF) module that addresses the often-overlooked challenge of fusing these features by anisotropically merging them into a unified representation. Extensive experiments demonstrate the advantages of FoundationSSC, achieving simultaneous improvements in both semantic and geometric metrics, surpassing prior bests by +0.23 mIoU and +2.03 IoU on SemanticKITTI. Additionally, we achieve state-of-the-art performance on SSCBench-KITTI-360, with 21.78 mIoU and 48.61 IoU.
Codehttps://github.com/D-Robotics-AI-Lab/FoundationSSC
@@ Introduction
Semantic Scene Completion (SSC) infers the complete geometry and semantics of a 3D scene from partial observations. It is essential for high-level 3D perception applications, such as intelligent driving (Zheng et al. 2024) and robotic navigation (Wu et al. 2025). Early SSC methods primarily utilised LiDAR input (Song et al. 2017; Roldao et al. 2020), but they faced challenges due to high hardware costs and limited resolution. As a result, there has been a recent shift towards camera-based SSC methods, which are gaining significant attention for their lower costs and richer texture information.
@@
In this paper, we propose FoundationSSC, a novel SSC framework designed to unleash the semantic and geometric priors from foundation models, as shown in Fig. 1 (b). Our approach operates on a core principle of dual decoupling. It begins with a unified foundation encoder that harnesses the power of VFMs to provide separate priors for the semantic and geometric branches, thereby addressing the trade-off at the feature source. These distinct priors are then processed through specialised, decoupled pathways aimed at maximising their potential. The resulting high-quality 2D representations are transformed into 3D space using a hybrid view transformation. Additionally, we introduce a novel Axis-Aware Fusion (AAF) module, which fuses the complementary 3D feature volumes in a way that respects their anisotropic characteristics, resulting in a unified 3D representation. Through these integrated designs, FoundationSSC effectively resolves the semantic-geometric coupling issues that have limited previous work, producing high-quality 2D and 3D feature representations. Consequently, our framework achieves state-of-the-art performance on SemanticKITTI (Behley et al. 2019) and SSCBench-KITTI-360 (Liao et al. 2022) datasets, simultaneously enhancing both geometric (Intersection over Union, IoU) and semantic (mean IoU, mIoU) metrics. In summary, our contributions are as follows:
@@
MonoScene (Cao and De Charette 2022) is a pioneering effort that extends the SSC task to cameras by employing a Features Line of Sight Projection (FLoSP) for projecting 2D features into 3D space. TPVFormer (Huang et al. 2023) introduces an efficient tri-perceptive view representation, created from image features. VoxFormer (Li et al. 2023a) employs a sparse-to-dense strategy, which propagates image features to dense voxels through deformable attention mechanism. OccFormer (Zhang et al. 2023) builds upon the Lift-Splat-Shoot (LSS) paradigm to facilitate feature lifting using both local and global transformer pathways.
Subsequent works have proposed more specialised solutions, often leveraging temporal cues from multiple frames (Li et al. 2024a; Lee et al. 2025), utilising more accurate geometric information obtained from stereo matching (Li et al. 2024b; Xue et al. 2024; Yu et al. 2024), exploiting spatial biases (Wang and Tong 2024; Bae et al. 2025), and integrating richer semantic knowledge derived from instance-level features (Jiang et al. 2024) or vision-language models (Wang et al. 2025; Cui et al. 2025). However, despite the diversity of these approaches, they are often limited by a tightly coupled architecture, which results in a trade-off between semantic and geometric features.
Our proposed FoundationSSC disentangles semantic and geometric features at both the source and processing pathway levels, effectively resolving the trade-off that has constrained earlier work.
@@
  • Dense depth map $\mathbf{Z}$: The final disparity map predicted by FoundationStereo, which is then converted to a metric depth map using camera intrinsics. This is primarily used for generating voxel proposals in the view transformation and geometric prior for the semantic context branch.
@@
Disparity-to-Depth Volume Mapping.
@@
To restore spatial structure, $\mathbf{\tilde{V}}_{depth}$ is reshaped into 3D volumes $\mathbf{V}_{depth}\in\mathbb{R}^{D_{depth}\times H\times W}$. A refinement module then employs a shallow 3D CNN encoder $\phi(\cdot)$ to enhance spatial coherence and smoothness, followed by a softmax operation along the depth dimension to produce the probabilistic depth distribution, $\mathbf{D}\in\mathbb{R}^{D_{depth}\times H\times W}$:
@@
Hybrid View Transformation
@@
Axis-Aware Fusion.
@@
To train the FoundationSSC, we follow MonoScene (Cao and De Charette 2022), utilise affinity losses $\mathcal{L}_{scal}^{geo}$ and $\mathcal{L}_{scal}^{sem}$ to optimise the scene-wise and class-wise metrics, in conjunction with the cross-entropy loss $\mathcal{L}_{ce}$ weighted by class frequencies. Additionally, we employ two auxiliary losses: a 2D semantic segmentation loss $\mathcal{L}_{s}$ for feature enrichment and a depth loss $\mathcal{L}_{d}$ (Li et al. 2024e) for geometric supervision. The total loss is a weighted sum of these components:
@@
where $\lambda_{d}$ and $\lambda_{s}$ are the weights for the auxiliary losses. Following preliminary experiments, we set $\lambda_{d}=0.001$ and $\lambda_{s}=1$, respectively.
@@
Ablation Study for Foundation Encoder.
@@
Ablation Study for DDVM.
We validate our DDVM module against various depth generation strategies in Table 5. The CGFormer (Yu et al. 2024) depth refinement module, which fuses monocular distribution with a pre-processed stereo depth map using neighbour attention, achieves 19.83 mIoU and 47.87 IoU. A key limitation of such a method is its reliance on information-lossy depth maps, which discard valuable probabilistic cues from stereo matching. We then explore the direct use of the disparity cost volume. A naive approach using analytical resampling (AR) degrades performance to 19.59 mIoU, confirming that merely accessing the cost volume is insufficient. In contrast, our DDVM, by replacing rigid resampling with a learnable mapping, achieves a state-of-the-art mIoU of 20.36. The performance gain demonstrates DDVM’s superior ability to preserve and leverage fine-grained geometric cues for SSC.
@@
Ablation Study for AAF.
Table 6 compares the AAF module against alternative 3D feature fusion strategies. We establish a baseline by using only the refined volume $\mathbf{F}_{vt}$, mirroring prior work, which yields 19.56 mIoU. A fusion approach using a vanilla 3D channel attention (3D-CA) module improves performance to 20.08 mIoU. However, the effectiveness of 3D-CA is constrained by an isotropic global pooling, a design that neglects directional cues critical for 3D understanding. In contrast, our AAF module is explicitly designed to be anisotropic. The module’s axis-specific fusion mechanism achieves the highest mIoU of 20.36. The substantial performance gain over the 3D-CA provides compelling evidence that modelling the anisotropic nature of 3D scenes is crucial for effective feature fusion.
@@ Conclusion
Appendix/Supplemental Material
In the appendix, we mainly provide implementation details and more experiment results.
Experimental Setup
Datasets. We evaluate our FoundationSSC on two widely-used datasets: SemanticKITTI (Behley et al. 2019) and SSCBench-KITTI-360 (Li et al. 2024d). SemanticKITTI provides RGB images of size $1226\times 370$ and includes 20 classes (19 semantic classes and one free class). The dataset comprises 10 training sequences, one validation sequence, and 11 testing sequences. SSCBench-KITTI-360 (Li et al. 2024d) offers input RGB images at $1408\times 376$ resolution. It contains 19 classes (18 semantic classes and one free class) with seven training sequences, one validation sequence, and one testing sequence. All evaluations on these two benchmarks focus on a defined spatial area: $51.2m$ in front of the vehicle, $25.6m$ to both the left and right, and $6.4m$ above the vehicle. Voxelizing this region results in a set of 3D voxel grids with a resolution of $256\times 256\times 32$, where each voxel measures $0.2m$ on each side.
Metrics. Following previous methods (Cao and De Charette 2022; Li et al. 2023a; Yu et al. 2024), we report the Intersection over Union (IoU) and mean IoU (mIoU) metrics for occupied voxel completion and voxel-wise semantic completion evaluation, respectively. These two metrics offer a comprehensive view of the model’s ability to reconstruct the scene in terms of both geometry and semantics.
Implementation Details
Image Encoder. We adopt the frozen FoundationStereo (Wen et al. 2025) as the image encoder and leverage its backbone, DepthAnythingV2 (Yang et al. 2024b), to extract multi-level image features. Since a ViT-based backbone is non-hierarchical, as in (Li et al. 2022), we utilise the last and strongest feature map to produce multi-scale feature maps. We then align them in both spatial and channel dimensions, and finally concatenate them along the channel dimensions into a stacked feature tensor. Note that DepthAnythingV2 originally outputs feature maps at a $1/14$ resolution. For easy access to subsequent feature utilisation, we interpolate the multi-level feature maps to $1/16$. After passing them through the FPN and stacking the multi-scale feature maps, the final image feature tensor $\mathbf{F}^{2D}$ is at a $1/8$ resolution.
The input resolution is $384\times 1280$ on SemanticKITTI and $384\times 1408$ on SSCBench-KITTI-360 to match the patch size of the FoundationStereo encoder.
Training Setup. We train FoundationSSC for 25 epochs on 4 NVIDIA 4090 GPUs, with a batch size of 4. It approximately consumes 22 GB of GPU memory on each GPU during the training phase. We employ the AdamW (Loshchilov and Hutter 2018) optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.99$ and set the maximum learning rate to $3\times 10^{-4}$. The cosine annealing learning rate strategy is adopted for the learning rate decay, where the cosine warmup strategy is applied for the first $5\%$ iterations.
Quantitative Results
To provide a more thorough comparison, we report additional quantitative results on SemanticKITTI validation set in Table 7. The state-of-the-art performance further validates the effectiveness of our approach in advancing 3D scene perception.
Qualitative Results
We report additional visualization results on SemanticKITTI validation set in Fig. 7.
Figure 6: Failure Cases.
Figure 7: More qualitative comparison results on SemanticKITTI validation set.
Method IoU mIoU
$\blacksquare$ road (15.30%)
$\blacksquare$ sidewalk (11.13%)
$\blacksquare$ parking (1.12%)
$\blacksquare$ other-grnd. (0.56%)
$\blacksquare$ building (14.1%)
$\blacksquare$ car (3.92%)
$\blacksquare$ truck (0.16%)
$\blacksquare$ bicycle (0.03%)
$\blacksquare$ motorcycle (0.03%)
$\blacksquare$ other-veh. (0.20%)
$\blacksquare$ vegetation (39.3%)
$\blacksquare$ trunk (0.51%)
$\blacksquare$ terrain (9.17%)
$\blacksquare$ person (0.07%)
$\blacksquare$ bicyclist (0.07%)
$\blacksquare$ motorcyclist (0.05%)
$\blacksquare$ fence (3.90%)
$\blacksquare$ pole (0.29%)
$\blacksquare$ traf.-sign (0.08%)
MonoScene 36.86 11.08 56.52 26.72 14.27 0.46 14.09 23.26 6.98 0.61 0.45 1.48 17.89 2.81 29.64 1.86 1.20 0.00 5.84 4.14 2.25
TPVFormer 35.61 11.36 56.50 25.87 20.60 0.85 13.88 23.81 8.08 0.36 0.05 4.35 16.92 2.26 30.38 0.51 0.89 0.00 5.94 3.14 1.52
OccFormer 36.50 13.46 58.85 26.88 19.61 0.31 14.40 25.09 25.53 0.81 1.19 8.52 19.63 3.93 32.62 2.78 2.82 0.00 5.61 4.26 2.86
VoxFormer 44.15 13.35 53.57 26.52 19.69 0.42 19.54 26.54 7.26 1.28 0.56 7.81 26.10 6.10 33.06 1.93 1.97 0.00 7.31 9.15 4.94
Symphonize 41.92 14.89 56.37 27.58 15.28 0.95 21.64 28.68 20.44 2.54 2.82 13.89 25.72 6.60 30.87 3.52 2.24 0.00 8.40 9.57 5.76
H2GFormer 44.69 14.29 57.00 29.37 21.74 0.34 20.51 28.21 6.80 0.95 0.91 9.32 27.44 7.80 36.26 1.15 0.10 0.00 7.98 9.88 5.81
CGFormer 45.99 16.87 65.51 32.31 20.82 0.16 23.52 34.32 19.44 4.61 2.71 7.67 26.93 8.83 39.54 2.38 4.08 0.00 9.20 10.67 7.84
HTCL 45.51 17.13 63.70 32.48 23.27 0.14 24.13 34.30 20.72 3.99 2.80 11.99 26.96 8.79 37.73 2.56 2.30 0.00 11.22 11.49 6.95
VLScene 44.69 17.83 63.10 31.10 24.40 0.20 24.90 33.40 30.70 1.80 3.60 18.30 26.00 8.10 35.30 4.30 2.60 0.00 12.10 11.90 6.30
Ours 47.91 20.36 67.37 34.12 21.70 0.77 27.96 37.08 41.20 5.94 7.07 20.12 28.61 10.95 41.06 5.02 4.15 0.04 13.01 12.65 8.01
Table 7: Quantitative results on SemanticKITTI validation set. Bold denotes the best performance.
Foundation Encoder IoU mIoU Params (M) Trainable Params (M) Training Memory (MB)
FoundationStereo-L 48.12 19.32 444.48 67.38 22030
FoundationStereo-S 47.52 18.05 126.04 61.12 21400
Table 8: Efficiency and performance comparison of FoundationSSC with Large vs. Small FoundationStereo encoders on SemanticKITTI hidden test set.
Failure Cases
Fig. 6 presents two challenging scenes that highlight the limitations of FoundationSSC in occluded regions. In the top example, the model fails to infer the presence of an intersection ahead, resulting in incomplete road structure prediction. In the bottom example, one vehicle is partially occluded by another. Due to the lack of instance-level supervision during training, the model struggles to delineate the boundaries of the occluded vehicle, leading to merged voxels between the two cars.
Results with Lightweight Foundation Encoder
To assess the efficiency of our method, we evaluate FoundationSSC using a lightweight backbone by replacing the FoundationStereo-L encoder with FoundationStereo-S. As shown in Table 8, this substitution significantly reduces computational cost while maintaining competitive performance on SemanticKITTI hidden test set.
Limitations
FoundationSSC currently operates on single-frame inputs. Although it achieves state-of-the-art performance, it does not leverage temporal information, which could help address cross-frame inconsistencies and enhance overall scene understanding. Incorporating temporal cues will be an important direction for our future work.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (62473276, 62573309), in part by the Natural Science Foundation of Jiangsu Province (BK20241918), and in part by the Research Fund of Horizon Robotics (H230666).